20 tablets weighing 2g was powdered. Exactly 50mg was transferred to

100mL volumetric flask and diluted with water. 10mL aliquot was

transferred to another 100mL and diluted to volume. The A value of the

solution was 0.652 and the absorptivity value was 15mg/mL/cm. What is

the mg of paracetamol per tablet from the following data? Please

provide a step by step solution.


1
Expert's answer
2021-02-17T12:41:52-0500

According to the absorption formula:

A = a b c,

where A - absorbance (0.652), a - absorptivity (15 mg/mL/cm), b - length of path (1 cm), c - concentration.

From here:

c = A / (a b) = 0.652 / (15 mg/mL/cm × 1 cm) = 0.0435 mg/mL

As a result, the second 100 ml flask contained 0.0435 mg/mL × 100 mL = 4.35 mg of paracetamol. From here, the first 100 mL flask contained 4.35 mg × 10 = 43.5 mg of paracetamol. It means that 50 mg of tablet contains 43.5 mg of paracetamol.

Since 20 tablets weigh 2 g, one tablet weighs 2 g / 20 = 0.1 g = 100 mg.

Finally, as 50 mg of tablet contains 43.5 mg of paracetamol, one tablet (100 mg) contains 43.5 mg × 2 = 87 mg of paracetamol.


Answer: 87 mg